40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 16 03:01:57 2015, ericq, Update, LSC, Old beatbox hooked back up beat_comparison.pngaLIGO_vs_beatbox.xml.zip
    Reply  Thu Jul 23 03:09:05 2015, ericq, Update, LSC, aLIGO demod board lives! beatbox_vs_aLIGO.png
       Reply  Thu Jul 30 14:42:03 2015, ericq, Update, LSC, aLIGO demod board lives! partialALSbudget.pngdemodDriveLevels.pngALScoherences.pngpartialALSbudget.zip
          Reply  Thu Jul 30 15:58:00 2015, ericq, Update, LSC, Beat note Alignment fluctuation effects measured RF_RIN.pngRF_RINspec.pngRFampCoh.xml.zip
             Reply  Tue Aug 4 03:02:30 2015, ericq, Update, LSC, Beat note Alignment fluctuation effects measured coherences.pnglinX.png
                Reply  Tue Aug 11 02:40:29 2015, ericq, Update, LSC, 50m delay lines - Rough calibrations 
Message ID: 11468     Entry time: Thu Jul 30 14:42:03 2015     In reply to: 11438     Reply to this: 11470
Author: ericq 
Type: Update 
Category: LSC 
Subject: aLIGO demod board lives! 

ALS is not currently limited by the demod board or whitening electronics.

The noise budget in the green locking paper shows the main noise sources to be these two, plus the residual fluctuations of the green PDH loop. 

So, one next step is AUX PDH noise budget. 

However, I wonder how much of the low frequency noise can be explained by instability of the beat alignement on the PSL table, and how this might be quantified. 


Yesterday, I put together a few measurements to asses whether the new demod board has moved us in the right direction. Specifically I measured the output of the phase tracker in the following states, adjusting the phase tracker gain to maintain a ~2kHz UGH (but no boost on):

  • Whitening chassis inputs terminated. BEAT_I input channels were given a 3000 count offset to give the phase tracker something to work with. This is a typical beatnote amplitude with the new RF amplifiers. 
  • aLIGO LSC demod board driven with an SRS SD345 at 30MHz. (First with +3dBm into the splitter, which is about what it sees with the green beatnotes, then with +13dBm into the splitter, to give the board the +10dBm LO it expects)
  • Arms locked with POX, POY. AUX laser temperature servos on. Green beatnotes in the 20-40MHz range. 

Results: The beat frequency spectrum is above the measured demod board and whitening chassis/ADC noise at all frequencies. It's a little close at 10Hz. 

One nice feature is that the beat spectra are far more similar to each other than they used to be. RMS noise is in the 300-400Hz range, which isn't mindblowing, but not terrible. On the order of 50 pm for each arm. Most of this comes from below 10Hz. 

Another thing to note is that, when we switch in the 50m cables, we should win a fair bit of Hz/V gain and push down these noises futher. (We're currently using 30m cables.)


By looking at some coherences, we can attribute some of the noise when IR locked to both colors of PDH loops. 

Specifically, the coherence with the Green PDH error implicates the residual frequency noise of the AUX laser above a few hundred Hz, whereas the feature from 20-50Hz is probably real cavity motion, not ALS sensing noise. Some of the 1-3Hz noise is from real suspension/stack resonances too. 


If it turns out that we do want to push the demod board noise down further, we could think about increasing the RF amplification. Driving the board harder translates directly to better noise performance. The 60Hz harmonics aren't so exciting, but not the end of the world.


Data files are attached, if you're in to that sort of thing. 

Attachment 1: partialALSbudget.png  291 kB  | Hide | Hide all
partialALSbudget.png
Attachment 2: demodDriveLevels.png  212 kB  | Hide | Hide all
demodDriveLevels.png
Attachment 3: ALScoherences.png  238 kB  | Hide | Hide all
ALScoherences.png
Attachment 4: partialALSbudget.zip  4.305 MB
ELOG V3.1.3-